Equal volumes, 50.0 ml of 0.54897 M solution of HCl and 0.9573 M NaOH solutions having an initial temperature of 18.5 degree Celsius react in a coffee cup calorimeter. The resultant solution records a temperature of 22.9 degree Celsius. since the solutions are mostly water, the solutions are assumed to have a density of 1.0g/ml and a specific heat of 4.184 j/g degree Celsius. The heat capacity of the calorimeter is 1.0*10^1 j/degree. Which reagent is limiting
1
Expert's answer
2021-08-22T07:29:53-0400

Number of moles of HCl and NaOH added


= 1000


MV



= 1/1000×50



=0.05


Mass of mixture =V×d=100×=100g


Heat evolved, q=msΔT=100×4.18×(27.5−21.0)


=100×4.18×6.5J=2717J=2.717 kJ




(b) The involved reaction is:


HCl(aq.)+NaOH(aq.)→NaCl(aq.)+H2O


ΔH= Heat evolved per mol


= 0.05−2.717